If in triangle ABC, AD is the altitude from A. given b>c, angle C =23degree & AD =abc / b^2-c^2.


then angle B equal to?


Grade:

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Nehal,

Ans:- AD=c sinB=abc/(b² - c²)

             or sinB=ab/(b² - c²)

              or sinB=sinA sinB/sin(B+C) sin(B-C)

                or sinB=sinAsinB/sinAsin(B-C)

                or sin(B-C)=1=sin90°

                     or B-C=90°

                        Hence B=113°(ans)

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Nehal !!!

 


Regards,

Askiitians Experts

SOUMYAJIT IIT_KHARAGPUR